# Facilitating Automated Online Consensus Building through Parallel Thinking

### v1

@@ 4.2. Results
Our dataset consists of 32 discussions, 2 by each of the 16 groups, with each group consisting of 3 human participants. In total these discussions consist of 16,656 words in 1,669 posts, of which the facilitator contributed 3,459 words in 217 posts. Each post has a corresponding timestamp and, for the LLM-generated posts, which hat was used to generate the text. Analysis of the discussion transcripts reveals that while there are promising indications of the efficacy of PTFA-based consensus building, there are also challenges in utilizing an LLM to make timely, appropriate contributions that consistently advance the discussion in a productive manner. The post-discussion surveys support this observation. The survey responses indicate that participants had more mixed feelings about the performance of the PTFA facilitator on a number of discussion attributes, including the extent to which the facilitator helped consensus decision making (see Figure 2(b)), when compared to a baseline facilitation approach. Below, we first present the survey results that provide an overview of participant experiences. We then examine the types of constructive conversation pattern observed in the discussions before analyzing the problematic timing and formulation of comments that likely contributed to the lower satisfaction scores.
4.2.1. Survey Responses
@@
Generally, the user experience with either facilitation model has been positive in terms of user experience (see Figure 2(a)) and reached consensus (see Figure 2(c)). However, despite positive contributions by the PTFA-based facilitator (Facilitation Model 1) towards being an effective and practical consensus building facilitator (see Section 4.2.2) negative effects (see Section 4.2.3) lead to mixed facilitator ratings. We observe this by the response to the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ to which the participation could indicate their agreement with 7 the responses: ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’. While participants react positively to Facilitation Model 0, as Figure 2(b) shows, the PTFA-based facilitator is rated much more diversely in the responses while still being positive to neutral overall.
The mixed performance of the PTFA facilitator also affects the user experience and their agreement with the consensus. As Figure 2(a) highlights, while the user experience is still positive, it is slightly dampened by some of the problematic LLM patterns (see Section 4.2.3). The user experience assessment is based on the question ‘How would you rate the user experience of the platform?’ to which the participation could indicate their satisfaction with 7 the responses: ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’. Moreover, the PTFA facilitator also dampens the agreement with the consensus, especially due to disrupting the consensus forming with the LLM’s difficulty to keep to specified phases (see Problematic Facilitation Patterns in Section 4.2.3). This consensus agreement assessment is based on the question ‘Do you agree with the consensus reached in this discussion?’ to which the participants could indicate their agreement with the same 7 responses as above.
@@
Figure 2. Survey responses for the questions around the rating of the facilitator models and the agreement with the resulting consensus. The legend in Figure 2(e) shows the possible responses used for the questions in Figure 2(b) and Figure 2(c). The 50% line divides the bars at 50% of participants.
[…]
Bar graph showing the share of survey responses for the question ‘How would you rate the user experience of the platform?’ rated into the 7 possible responses ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’ as well as bar graphs showing the share of survey responses for the questions ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ and ‘Do you agree with the consensus reached in this discussion?’ rated into the 7 possible responses ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’.
4.2.2. Constructive Facilitation Patterns
@@
We observed several instances in which the PTFA facilitator made productive contributions to the discussion that are distinct from the types of contributions provided by the human participants. We identify four types of useful intervention, each of which was contributed by a particular thinking hat. In the following, we describe these useful interventions. Example discussion excerpts of positive contributions are presented in the appendix.
@@
Highlighting and exploring emotional responses of participants (Red Hat). The human participants mostly gave shallow opinions about suggestions; they typically said little more than whether or not they like an idea. By asking focused questions about the participants’ emotional reactions and feelings, the PTFA facilitator helped them make better informed decisions about their existing ideas and generate new ones better aligned with their preferences.
@@
Identifying positive/negative qualities of suggestions that meet the participants needs (Yellow/Black Hat) Identifying strengths and weaknesses of suggestions, the PTFA facilitator provided a deeper analysis of proposals than that typically provided by the human participants independently. These comments highlighted particular features of ideas that aligned or conflicted with participants’ expressed interests. Contributions of this type usually expanded on a positive or negative comment made by a participant.
@@
Providing factual information about the discussion topic (White Hat) The LLM is able to generate relevant facts learned from its training data. It did this during discussions either when participants expressed that they did not know some factual information (for example, the runtime of a film) or explicitly asked the facilitator for such information. While it is possible for LLMs to hallucinate details, almost all the facts it generated were correct. One exception was a comment that a theater production would open in the future, even though it had actually already started. Inaccuracies like this are due to the cut-off date of the data used to train the model rather than hallucination of facts.
@@
Introducing new suggestions to the discussion (Green Hat) The PTFA facilitator regularly provided or extended suggestions. These suggestions were often well received by participants and then selected as the consensus decision. The LLM has a tendency to combine or extend prior suggestions - particularly when generating ideas for a shared activity. While these extensions were also sometimes well received, on other occasions the participants joked about how many components the facilitator was adding to the activities.
4.2.3. Problematic Facilitation Patterns
@@
While the PTFA facilitator is capable of making useful contributions, these are offset by the many comments that were unhelpful, or even counterproductive. We observed two main categories of issues: poor management of the discussion phase and inconsistent response timing. Example discussion excerpts of problematic contributions are presented in the appendix.
Poor Phase Management
@@
The most significant of these issues is in scheduling the phases of the consensus building discussion. The LLM is prompted to facilitate a divergent first discussion stage encouraging idea generation, followed by a convergent second stage of consensus formation. In practice, the facilitator stayed in the first stage throughout the discussions, with the green and blue hat personas consistently asking for further suggestions until the end of the 20 minute period, including when participants were trying to form a consensus or explicitly stated that they had reached a consensus they were happy with. It also asked the participants to generate new suggestions while they were already in the process of sharing new suggestions. Often, these inappropriate requests for more ideas were ignored by the participants. However, in one discussion a consensus was reached until the facilitator intervened with more suggestions. The new ideas split the participants’ preferences without enough time for them to discuss and agree on a new unanimous decision.
Hence, to build an effective facilitator based on LLMs, it is thus necessary to allow the facilitator to be aware of phases and times and fine-tune the facilitator to judge the state of the discussion from the participants’ perspective.
Inconsistent response timing
@@
Another issue is that the PTFA facilitator sometimes did not respond for long periods of time. In one instance, it makes its first contribution 8 minutes into the session, with a comment “starting” the discussion - despite the participants already having generated many ideas. Similarly, it sometimes ignores questions explicitly asked of it. On the other hand, the PTFA facilitator may intervene too much - responding after every individual comment without allowing the participants to develop their ideas. These repeated messages are typically those asking for more suggestions, as described above.
Hence, it is necessary to be aware that LLMs can interact differently in seemingly similar situations and fine-tune the LLM’s intervention rate based on the state of the discussion, with respect of the content, and its previous interventions.
4.3. Discussion
@@
This study provides insight in to both the potential and limitations of LLM-augmented discussions. Analysis of the 32 facilitated discussions shows that in a PTFA-based framework, LLMs can make valuable contributions that are otherwise neglected by human participants — particularly through idea generation, emotional probing, and deeper analysis of positive and negative qualities of ideas. These findings demonstrate the potential utility of LLMs in discussion facilitation, offering examples of how a PTFA could automate skilled facilitation to improve outcomes of consensus building. Despite this, the survey results reveal that users had more mixed opinions of the PTFA facilitator when compared to a well-received baseline facilitator on several discussion attributes. This is supported by the qualitative analysis of discussion transcripts, in which we identified many instances of the PTFA facilitator making unhelpful or inappropriate contributions to the discussion.
@@
5. Conclusions
This paper explored the potential of AI agents in supporting consensus building processes in a natural conversation scenario. We introduced the Parallel Thinking-based Facilitation Agent (PTFA) that utilized parallel thinking-based strategies aligned with general-purpose LLMs. A real-world pilot user study, including 48 participants, was conducted, summarizing the PTFA’s ability along with the challenging issues. We not only highlight insights how LLMs are suitable but also noted some limitations and open challenges, thus providing directions for utilizing LLMs effectively for facilitation agents. Moreover, based on the user study, we constructed a novel dataset rich in conversational content among participants and between participants and PTFA. Future work will focus on addressing the open challenges identified in the user study.
Acknowledgements.

### v2

@@ 4.2 Results
Analysis of the discussion transcripts reveals that while there are promising indications of the efficacy of PTFA-based consensus building, there are also challenges in utilizing an LLM to make timely, appropriate contributions that consistently advance the discussion in a productive manner. The post-discussion surveys support this observation. The survey responses indicate that participants had more mixed feelings about the performance of the PTFA facilitator on a number of discussion attributes, including the extent to which the facilitator helped consensus decision making (see Figure 4(b)), when compared to a baseline facilitation approach. Below, we first present the survey results that provide an overview of participant experiences. We then examine the types of constructive conversation patterns observed in the discussions before analyzing the problematic timing and formulation of comments that likely contributed to the lower satisfaction scores.
4.2.1 Survey Responses
@@
Generally, the user experience with either facilitation model has been positive in terms of user experience (see Figure 4(a)) and reached consensus (see Figure 4(c)). However, despite positive contributions by the PTFA-based facilitator (Facilitation Model 1) towards being an effective and practical consensus building facilitator (see Section 4.2.2), negative effects (see Section 4.2.3) lead to mixed facilitator ratings. We observe this in the response to the question ‘How would you rate the extent to which the facilitator in the discussion helped consensus decision-making?’ to which the participant could indicate their agreement with one of the statements: ‘Strongly Agree’, ‘Agree’, ‘Somewhat Agree’, ‘Neutral’, ‘Somewhat Disagree’, ‘Disagree’ and ‘Strongly Disagree’. While participants reacted more positively to the PTFA-based facilitator - being slightly above 50% - in comparison to Facilitation Model 0 - being slightly below 50% - as Figure 4(b) shows, overall, the rating of both is mixed.
The mixed performance of the PTFA facilitator also affects the user experience and their agreement with the consensus. As Figure 4(a) highlights, while the user experience is still positive, it is slightly dampened by some of the problematic LLM patterns (see Section 4.2.3). The user experience assessment is based on the question ‘How would you rate the user experience of the platform?’ to which the participation could indicate their satisfaction with 7 the responses: ‘Very Satisfied’, ‘Satisfied’, ‘Somewhat Satisfied’, ‘Neutral’, ‘Somewhat Unsatisfied’, ‘Unsatisfied’ and ‘Very Unsatisfied’. Moreover, the PTFA facilitator also slightly dampens the agreement with the consensus, especially due to disrupting the consensus forming with the LLM’s difficulty to keep to specified phases (see 7 in Section 4.2.3). This consensus agreement assessment is based on the question ‘Do you agree with the consensus reached in this discussion?’ to which the participants could indicate their agreement with the same 7 responses as above.
@@
Figure 4: Survey responses for the questions around the rating of the facilitator models and the agreement with the resulting consensus. The legend in Figure 4(e) shows the possible responses used for the questions in Figure 4(b) and Figure 4(c). The 50% line divides the bars at 50% of participants.
4.2.2 Constructive Facilitation Patterns
@@
We observed cases in which the PTFA facilitator made productive contributions to the discussion that were distinct from the types of contributions provided by the human participants. We identify four types of useful intervention, each of which was contributed by a particular thinking hat. In the following, we describe these useful interventions. Example discussion excerpts of positive contributions are presented in Figure 5 and Figure 6.
(a) The red hat externalizing its analysis of the participants comment.
(b) The white hat responding with accurate factual information to a question posed by a participant.
Figure 5: Examples of constructive facilitator contributions
(a) The yellow hat highlighting positives of a movie. The white hat providing accurate factual information about a movie in response to a participant expressing their unfamiliarity with it.
(b) The yellow hat highlighting positive aspects of ideas that resonate with the participants.
Figure 6: Examples of constructive facilitator contributions
@@
Highlighting and exploring emotional responses of participants (Red Hat). The human participants mostly gave shallow opinions about suggestions; they typically said little more than whether or not they liked an idea. By asking focused questions about the participants’ emotional reactions and feelings, the PTFA facilitator helped them make better informed decisions about their existing ideas and generate new ones better aligned with their preferences.
@@
Identifying positive/negative qualities of suggestions that meet the participants needs (Yellow/Black Hat). Identifying strengths and weaknesses of suggestions, the PTFA facilitator provided a deeper analysis of proposals than that typically provided by the human participants independently. These comments highlighted particular features of ideas that aligned or conflicted with participants’ expressed interests. Contributions of this type usually expanded on a positive or negative comment made by a participant.
@@
Providing factual information about the discussion topic (White Hat). The LLM is able to generate relevant facts learned from its training data. It did this during discussions, either when participants expressed that they did not know some factual information (for example, the runtime of a film) or explicitly asked the facilitator for such information. While it is possible for LLMs to hallucinate details, almost all the facts it generated were correct. One exception was a comment that a theater production would open in the future, even though it had actually already started. Inaccuracies like this are due to the cut-off date of the data used to train the model rather than the hallucination of facts.
@@
Introducing new suggestions to the discussion (Green Hat). The PTFA facilitator regularly provided or extended suggestions. These suggestions were often well-received by participants and then selected as the consensus decision. The LLM has a tendency to combine or extend prior suggestions, particularly when generating ideas for a shared activity. While these extensions were also sometimes well received, on other occasions, the participants joked about how many components the facilitator was adding to the activities.
4.2.3 Problematic Facilitation Patterns
@@
While the PTFA facilitator is capable of making useful contributions, these are offset by the many comments that were unhelpful or even counterproductive. We observed two main categories of issues: poor management of the discussion phase and inconsistent response timing. Example discussion excerpts of problematic contributions are presented in Figure 7.
(a) Asking for idea suggestions repeatedly and at an inappropriate time. These comments begin 4 minutes before the end of the 20 minute discussion period.
(b) The red hat externalizing its analysis of a participant’s comment.
Figure 7: Examples of problematic facilitator behaviors.
@@
Poor Phase Management. The most significant of these issues is in scheduling the phases of the consensus building discussion. The LLM is prompted to facilitate a divergent first discussion stage, encouraging idea generation, followed by a convergent second stage of consensus formation. In practice, the facilitator stayed in the first stage throughout the discussions, with the green and blue hat personas consistently asking for further suggestions until the end of the 20-minute period, including when participants were trying to form a consensus or explicitly stated that they had reached a consensus they were happy with. It also asked the participants to generate new suggestions while they were already in the process of sharing new suggestions. Often, these inappropriate requests for more ideas were ignored by the participants. However, in one discussion, a consensus was reached until the facilitator intervened with more suggestions. The new ideas split the participants’ preferences without enough time for them to discuss and agree on a new unanimous decision. Hence, to build an effective facilitator based on LLMs, it is necessary to allow the facilitator to be aware of phases and times and fine-tune the facilitator to judge the state of the discussion from the participants’ perspective.
@@
Inconsistent response timing. Another issue is that the PTFA facilitator sometimes did not respond for long periods of time. In one instance, it makes its first contribution 8 minutes into the session, with a comment “starting” the discussion - despite the participants already having generated many ideas. Similarly, it sometimes ignores questions explicitly asked of it. On the other hand, the PTFA facilitator may intervene too much, responding after every individual comment without allowing the participants to develop their ideas. These repeated messages are typically those asking for more suggestions, as described above. Hence, it is necessary to be aware that LLMs can interact differently in seemingly similar situations and fine-tune the LLM’s intervention rate based on the state of the discussion, with respect to the content, and its previous interventions.
4.3 Discussion
@@
This study provides insight into both the potential and limitations of LLM-augmented discussions. Analysis of the 32 facilitated discussions shows that in a PTFA-based framework, LLMs can make valuable contributions that are otherwise neglected by human participants, particularly through idea generation, emotional probing, and deeper analysis of positive and negative qualities of ideas. These findings demonstrate the potential utility of LLMs in discussion facilitation, offering examples of how a PTFA could automate skilled facilitation to improve outcomes of consensus building. Despite this, the survey results reveal that users had more mixed opinions of the PTFA facilitator when compared to a well-received baseline facilitator on several discussion attributes. This is supported by the qualitative analysis of discussion transcripts, in which we identified many instances of the PTFA facilitator making unhelpful or inappropriate contributions to the discussion.
@@
5 Conclusions
This paper explored the potential of AI agents in supporting consensus building processes in a natural conversation scenario. We introduced the Parallel Thinking-based Facilitation Agent (PTFA) that utilized parallel thinking-based strategies aligned with general-purpose LLMs. A real-world pilot user study, including 48 participants, was conducted, demonstrating the PTFA’s ability while also identifying a number of remaining challenges. We not only highlight insights on how LLMs are suitable but also note some limitations and open challenges, thus providing directions for utilizing LLMs effectively for facilitation agents. Moreover, based on the user study, we constructed a novel dataset rich in conversational content among participants and between participants and PTFA. Future work will focus on addressing the open challenges identified in the user study.
{credits}
5.0.1 Acknowledgements
This work was supported by JST CREST JPMJCR20D1, and JSPS KAKENHI Grant Number JP24K20900. Additional support for this work was provided by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (EP/V022067/1) on Citizen-Centric AI Systems (https://ccais.ac.uk/) and through the FEVER Programme Grant (EP/W005883/1).